OpenStudy (anonymous):

Verify cos(a+b)/cos(a-b)=1-tan(a)tan(b)/1+tan(a)tan(b)

OpenStudy (anonymous):

See if you follow this. reference line numbers for explanations, if necessary.

OpenStudy (anonymous):

Line 1: sum/diff identities for cosine Line 2: mutltiply by 1 in the form of the reciprocal of cos(a)cos(b) over itself; this give the ! in numerator and denominator Line 3: cancellation property of algebra Line 4: Quotient property tan(a) = sin(a)/cos(a)

OpenStudy (anonymous):

@gryphon i don't know how do this... can you start it a bit for me? This is extremely difficult.

OpenStudy (anonymous):

I just did it (see the pdf in my first post-- did you not see it?); and I explained it line by line in the second post.

OpenStudy (anonymous):

oooh sorry, i've been at this the entire day haha

OpenStudy (anonymous):

thanks a lot, you're saving me here

OpenStudy (anonymous):

do you need any more explanation of the line by line?

OpenStudy (anonymous):

The key to this is line 2 where you force the required terms to become 1.

OpenStudy (anonymous):

ah ok, i didn't understand how to get to that second line

OpenStudy (anonymous):

Also going from line 1 to line 2, you have to commute the terms of the denominator

OpenStudy (anonymous):

I have this other problem, it's 2tan(x/2) = sin^2x+1-cos^2x/sinx(1+cosx)

OpenStudy (anonymous):

and again thank you very much, i've had to do like 15 of these and they're all incredibly hard but these 2 have been impossible

OpenStudy (anonymous):

np, please double check that you entered the identity correctly before I try. thanks

OpenStudy (anonymous):

It's 2tan(x/2) = sin^2 x+1-cos^2 x / sinx(1+cosx) That's what it says on the paper.

OpenStudy (anonymous):

ok, i think i have something close... give me a minute

OpenStudy (anonymous):

ok, I got it, give me a second to clean it up and post it

OpenStudy (anonymous):

wow thanks, i really appreciate you taking your time to do this.

OpenStudy (anonymous):

np! these aren't that hard once you get the hang of it. you just have to keep at it!

OpenStudy (anonymous):

u good?

OpenStudy (anonymous):

yep, i have 2 questions left that have to do with 0,2pi intervals. could you help me with that? No problem if you have to go you've done enough for me already.

OpenStudy (anonymous):

sorry, I would have liked to help, but I logged out before reading your last post.